A 2

                                                                                                                        http://staff.washington.edu/fox/tex/thesis.html

A

A1

A A A A 2

A3AA

A A A

A?

4 A

\chapter_{A}\TeX\\_{A}
  \chapter{A Brief\\Description of \TeX}
  The \TeX\ formatting program is the creation of
  Donald Knuth of Stanford University.

AL^{A}TEX User’s Guide?L^{A}TEX Companion? TEXbook

A^{5}?_{A}?_{A}? AA ?
’απoθανϵι̂νθϵ́λω

thesis.cls_{A}?

A

% LaTeX thesis control file
\documentclass[11pt,twoside]{thesis}
\begin{document}
% preliminary pages
%
\prelimpages
\include{prelim}
% text pages
%
\textpages
\include{chap1}
\include{chap2}
\include{chap3}
\include{chap4}
% bibliography
%
\bibliographystyle{plain}
\bibliography{thesis}
% appendices
%
\appendix
\include{appxa}
\include{appxb}
\include{vita}
\end{document}
Aoneside10pt11pt12pt
\includeonly?
\\\\
\appendix_{A}
\section\subsection?\\
Achapternotesendnotes
\printendnotes

A?

fullpage

?

leftfullpagefullpage
 \begin{figure}[p]% the left side caption
   \begin{leftfullpage}
     \caption{ . . . }
   \end{leftfullpage}
 \end{figure}
 \begin{figure}[p]% the right side space
   \begin{fullpage}
      . . .
      ( note.. no caption here )
   \end{fullpage}
 \end{figure}
 \begin{figure}[p]
    \begin{leftfullpage}
       \caption{ . . . }
    \end{leftfullpage}
 \end{figure}
 \begin{figure}[p]% the right side space
   \begin{xtrafullpage}
      . . .
      ( note.. no caption here )
   \end{xtrafullpage}
 \end{figure}
leftfullpagextrafullpage
xtrafullpagefullpage

AA

rotating
  \usepackage[figuresright]{rotating}
/usr/local/share/texmf/doc/...
 \begin{sidewaystable}
     ...
     \caption{ . . . }
 \end{sidewaystable}
\caption?
  \caption[toc]{text}
  \caption(-){(continued)}

(continued)

  \caption(a)[toc]{text}
  \caption(x){text}

x

A

\Title\Author\Program\Year\titlepage^{6}
  \Title{The Suitability of the \LaTeX\ Text Formatter\\
     for Thesis Preparation by Technical and\\
     Non-technical Degree Candidates}
  \Author{Jim Fox}
  \Program{Computing & Communications}
  \Year{1999}
  \titlepage
\Degreetext
\Degree{degree name}
\School{school name}
\Degreetext{degree text}
\titlepage
\Chair\Signature\signaturepage
  \Chair{Name of Chairperson}{title}{Chair's department}
  \Signature{Name of Committee member}
  \Signature{Name of Committee member}
  \Signature{etc}
  \signaturepage
\Chair
\Chair
\Chair\thesissignaturepage
\thesisquoteslip
\doctoralquoteslip
\doctoralabstractquoteslip

?

\quoteslip
  \quoteslip{Extensive copying . . . may be found.}
\abstract
  \abstract{This sample . . . `real' dissertation.}

A

\acknowledgments
  \acknowledgments{The author wishes . . . {\it il miglior fabbro}.\par}}
\dedication
\vita

A A A

\small\large\Large
  \usepackage{newcent}

A

http://staff.washington.edu/fox/tex/thesis.html
/usr/local/share/texmf/tex/latex/uw/

an egocentric imitation, actually See Appendix A to obtain the source to this thesis and the style file. See, for example, W. Shakespeare[?] for an early treatment. We mean the L^{A}TEX 2ε version of L^{A}TEX. Earlier versions, now called L^{A}TEX2.09 were much different. “Exactly” specifically excludes the inherent variety in print devices. Although many TEX-formatted documents contain no mathematics except the page numbers, it seems appropriate that this paper, which is in some sense about TEX, ought to demonstrate an equation or two. Here then, is a statement of the Nonsense Theorem. Assume a universe E and a symmetric function $ defined on E, such that for each $^{yy} there exists a $\overline{yy}, where $^{yy} = $\overline{yy}. For each element i of E define S(i) = ∑_{i} $^{yy} + $\overline{yy} + 0. Then if RR is that subset of E where 1 + 1 = 3, for each i

lim_{$→∞} ∫ Sdi = \{ \begin{matrix}0, & if i ∉ RR; \\ ∞, & if i ∈ RR. \end{matrix}

Actually, it wasn’t. It included a footnote—unusual for title pages.